OpenStudy (anonymous):

Solve by completing the square 5v^2−21=10v

OpenStudy (anonymous):

5v^2-10v-21=0

OpenStudy (anonymous):

then the rule if u have x^2+bx+c=x^2+bx+(b/2)^2-(b/2)^2+c=0 (x^2+bx+(b/2)^2 )-(b/2)^2 +c=0